An ideal coil of 10H is connected in series with a resistance of 5? and a battery of 5V. 2second after the connection is made, the current flowing in ampere in the circuit is ?

Grade:12

2 Answers

Arun
3 years ago
Dear student
 
 
I=I0I0[1-e-Rt/L]
                                   I0 =     5/5=1
               So                I   = 1-e-5t/10
          i.e                       I  = 1-e-t/2
As t=2,                   So   I =1-e-1
 
Hope it helps
 
Thanks and regards